OpenStudy (anonymous):

what is the factored form of 6x^2+8x+2

OpenStudy (anonymous):

Welcome to Open Study. Basic equation to solve. Start by solving 6x^2

OpenStudy (zzr0ck3r):

so we want two numbers that multiply to 6*2 = 12 but add to 8 6 and 2 are two such numbers so expand the middle term \(6x^2+8x+2\iff 6x^2+6x+2x+2\) and now factor what you can from the first two terms \(6x^2+6x+2x+2 = 6x(x+1)+2x+2\) now factor what you can from the last two terms to make them look like the first two terms \(6x(x+1)+2x+2=6x(x+1)+2(x+1)\) now factor out the \((x+1\) that both terms share \(6x(x+1)+2(x+1) = (x+1)(6x+2)= 6x^2+8x+2\)

OpenStudy (zzr0ck3r):

I am not sure what @vzfreakz is refering to when he/she says solve \(6x^2\). But this is how I go about doing these types of problems...

OpenStudy (anonymous):

Oh Okay! Thanks a lot, i'm not really fond with math so that really helped :)